ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

18. Teaching and Popularization of Mathematics

1991 MS Classification: 00A06 (General Mathematics) Baidyk, Tatyana N., International Renaissance Foundation P. O. Box 17, Kiev-30, 252030, Ukraine Evaluation of Mathematician Achievements International Science Education Program has experience in announcing and organization of competition for professors and academician scientists in mathematics. These competitions were began in Ukraine from 1994. The grants are paid due to financial support of G. Soros. Methods of rating evaluation is based on the achievements of scientists: number of the monographs and the text-books, aspirant number, lectures per year, number of conference papers, citation index and impact-factors of science magazines. We’ve supported more than 200 scientists from the different cities of Ukraine. New competition’ll be organized by ISEP together with Ministry of Science and Technology of Ukraine. We hope that support of scientist and their creative work is important for development of mathematics not only in Ukraine.